(NPV) A company has the choice between two different types of machines. Machine A costs less, but it also has a shorter life expectancy of two years. B costs more but lasts longer for four years. The expected cash flows after taxes for the two different types are as follows:

Machine

0

1

2

3

4

A

(10,000)

8,000

8,000

B

(12,000)

5,000

5,000

5,000

5,000

The cost of money of the firm is 10%.  Analyze the two options and advise the company which one is better.

show how to solve with formula and with financial calculator
Answer #1

To compare these two machines we should calculate NPV of both projects and then calculate Equivalent annual cash flows to compare projects with unequal lives.

NPV of Project A =PV of Cash Flows -Initial Investment =8000*(1-(1+10%)^-2)/10%-10000 =3884.2975
NPV of Project B =PV of Cash Flows -Initial Investment =5000*(1-(1+10%)^-4)/10%-12000 =3849.3272

Equivalent annual Cash flows of A using Financial calculator
N=2;I/Y =10%;PV =3884.2975;CPT PMT
PMT = 2238.10


Equivalent annual Cash flows of B using Financial calculator
N=4;I/Y =10%;PV =3849.3272;CPT PMT
PMT = 1214.35

Since PMT of Project A is higher it should be selected.